# Mixed Dynamics In Linear Networks: Unifying the Lazy and Active Regimes

### v1

@@ Appendix D The Gradient Flow Dynamics of $A_{\theta(t)}$ in Lazy Regime
Lemma D.1.
@@ E.1 Saddle-to-Saddle Regime
The goal of this section is to show that at the end of the first NTK regime, the singular vectors of $A_{\theta(t)}$ are roughly aligned with $A^{*}$. Moreover, the alignment remains to be good throughout the mixed regime. The following generalization of lemma B.6 and lemma B.7 will be useful.
@@
Using the same technique as B.6 we have
@@
Proceeding as in lemma B.7, we have
@@
We use induction lemma B.3 to prove the theorem. For the first requirement, notice that for every $i=1,2,\ldots,w$,
@@ E.2 First NTK Regime
The noise is clearly of order $O(\sigma^{2}\sqrt{wd})$. This completes the verification of $P_{1}$. By lemma B.9, $\|A_{\theta(T_{1}^{\prime})}-\Sigma\|_{op}\leq\sigma^{2}wO(\sigma^{2}w)d^{-\text{min}(\frac{\gamma_{w}-1}{8},\frac{1}{4})}$. By Lemma B.4 we are done.

### v2

@@ Appendix D The Gradient Flow Dynamics of $A_{\theta(t)}$ in Lazy Regime
In this section, we use a stability argument to show that if $\sigma^{2}w$ is infinitely larger than $d$, then the algorithm cannot converge.
Proposition D.1.
@@ Appendix E The Gradient Flow Dynamics of $A_{\theta(t)}$ in Active Regime
In section C, we proved stability for gradient flow dynamics for $A_{t}$. In this section, we prove similar stability statements for $A_{\theta(t)}$, using the approximation results from section B. We also summarize the behavior of $A_{\theta(t)}$ in section E.3.
@@ E.1 Saddle-to-Saddle Regime
The goal of this section is to show that at the end of the first NTK regime, the singular vectors of $A_{\theta(t)}$ are roughly aligned with $A^{*}$. Moreover, the alignment remains to be good throughout the mixed regime. The following generalization of lemma C.6 and lemma C.7 will be useful.
@@
Using the same technique as C.6 we have
@@
Proceeding as in lemma C.7, we have
@@
We use induction lemma C.3 to prove the theorem. For the first requirement, notice that for every $i=1,2,\ldots,w$,
@@ E.2 First NTK Regime
The noise is clearly of order $O(\sigma^{2}\sqrt{wd})$. This completes the verification of $P_{1}$. By lemma C.9, $\|A_{\theta(T_{1}^{\prime})}-\Sigma\|_{op}\leq\sigma^{2}wO(\sigma^{2}w)d^{-\text{min}(\frac{\gamma_{w}-1}{8},\frac{1}{4})}$. By Lemma C.4 we are done.
@@ Appendix F Gradient Descent Dynamics and Proof of Theorem 2
In this section we prove that gradient descent dynamics of $A_{\theta(t)}$ is well-approximated by gradient flow dynamics of $A_{\theta(t)}$.